OpenStudy (anonymous):

Consider the line y=-4x+5 . Find the equation of the line that is perpendicular to this line and passes through the point (-5,-6) . Find the equation of the line that is parallel to this line and passes through the point (-5,-6) .

OpenStudy (shamil98):

perpendicular slopes = inverse parallel slopes = the same Plug in the values to find b for the perpendicular line and do the same for the parallel line.

OpenStudy (shamil98):

For the parallel line you would for example : -6 = -5 (-4) + b -6 = 20 + b -26 = b so, y = mx + b y = -4x - 26 for the coordinates ( -5,-6)

OpenStudy (campbell_st):

oops.... well the perpendicular line has a slope that is the negative reciprocal of the slope belonging to the given line... not the inverse...